OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (anonymous):

(A) -4x - 9y = -19 -10y = -30 (B) 4x + 3y = 5 2y = -6 (C) 7x - 3y = -11 9x = -6 (D) 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

The solution is (-2, 3)

OpenStudy (anonymous):

I know that the choices cannot be B or C.

OpenStudy (anonymous):

Can u help me? @ganeshie8 @Hero @HelpBlahBlahBlah @pooja195

OpenStudy (anonymous):

D

OpenStudy (zale101):

Have you tried solving the system of linear equations for A and D?

OpenStudy (anonymous):

I think for those choices are (-2, 3). That's why I got stuck

OpenStudy (zale101):

For A -4x - 9y = -19 -10y = -30 I used substitution, where i solve for y for the equation -10y = -30 and got y=3, then plug y=3 for the equation -4x - 9(3) = -19 \(-4x-27=-19\) add 27 to both sides \(-4x=8\) divide -4 to both sides to solve for x, and you'll get x-2 so we got x=-2 and y=3